OpenStudy (anonymous):

In a recent year, suppose 24% of Americans identified themselves as liberal, and 28% as conservative. If 750 Americans are chosen at random, how many people who identified themselves as liberals and how many as conservatives would you expect?

OpenStudy (anonymous):

liberal:$$ 750 * 24 \% = 750 * 0.24 = 180 $$

OpenStudy (anonymous):

This is pretty straightforward. If you have a total of 750 Americans, you are looking for 28% of this. Which is 28% * 750 which can be rewritten as .28 * 750. So this amounts to an expected value of 210 conservative Americans.